Please help solve this question, with work shown if possible:
solve the following systems of linear equations using any algebraic method. If possible, check your solutions.
3y=11-2x
3x=y-11
Thank you!

Found 2 solutions by rothauserc, Boreal:
Answer by rothauserc(4718)

1) 2x + 3y = 11
:
2) 3x -y = -11
:
solve equation 1) for x
:
x = (-3y +11) / 2
:
now substitute for x in equation 2)
:
3( (-3y + 11 / 2) ) -y = -11
:
( -9y + 33 ) / 2 -y = -11
:
multiply both sides of = by 2
:
-9y + 33 -2y = -22
:
-11y = -55
:
***********************
y = 5
x = (-3(5)+11) / 2 = -2
***********************
:
check the equations in the problem
:
3y = 11-2x
3x = y-11
:
3(5) = 11 -2(-2)
15 = 15
:
3(-2) = 5-11
-6 = -6
:
our answer checks
:

Answer by Boreal(15235)

3y=11-2x
3x=y-11
rewrite
2x+3y=11
3x-y=-11
multiply the bottom by 3
9x-3y=-33
add to the top, and y disappears
11x=-22
x=-2
substitute into the second equation 3x-y=-11
-y=-5,
y=5
check in both of the top equations
15=11-2(-2), checks
-6=5-11 checks
(-2.5)